Determine how many solutions exist, and use either elimination or substitution to find the solutions (if any).

x + y = 3 and y = x + 3

Answer by mananth(16949)

x+y=3 .............1
-x+y=3 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-1 x -1 y = -3
-1 x 1 y = 3
Add the two equations
-2x =0
/-2
x= 0.00
plug value of x in (1)
x+y=3
y=3

unique solution ( 0,3)